# ViewActive: Active viewpoint optimization from a single image

### v4

@@ I INTRODUCTION
Humans can anticipate the best viewpoints for a given object from only a single observation, thanks to their extensive experience and observation of various objects. This benefits from humans’ ability of spatial visualization and mental rotation [1, 2] (see the left half of Fig. LABEL:fig:Banner), which allows them to proactively adjust their position to gather more information. The viewpoint’s quality significantly impacts a robot’s perception and interaction with the surroundings. Poorly chosen viewpoints, often referred to as “accidental views”[3, 4], can obscure critical features of an object. In contrast, “canonical views” [5] provide perspectives that reveal the most informative features of an object, significantly enhancing the robot’s performance in tasks such as object detection[6, 7, 8], navigation[9, 10], 3D reconstruction [11], and manipulation[12, 13]. Therefore, the ability to actively avoid accidental viewpoints and obtain better viewpoint is crucial for robots. This is known as active perception[14, 15], “where a robot purposefully selects what to perceive, how, and when to perceive it, driven by its current goals and interaction with the environment”.
@@ II Background & related works
The concept of “viewpoint” has been extensively discussed in computer graphics and virtual reality. The ability to select an optimal viewpoint is essential for accurately perceiving and interacting with objects in downstream tasks. In this section, we begin with the methods used to evaluate viewpoint quality and then discuss the challenges of selecting and optimizing viewpoints.
@@ II-A Viewpoint Quality Evaluation
Currently, we do not have a unified definition of a ”good viewpoint” [18, 19]. Since Palmer et al. first introduced the concept of a ”Canonical Perspective” [20] in 1981, a wealth of work has emerged on how to evaluate viewpoints and select the best viewpoint. Various existing viewpoint theories have evaluated viewpoints based on different criteria, such as area [21, 22], silhouette [23, 24], depth [25, 26], surface curvature [27, 28], semantics [29, 30], aesthetics[31], etc.

### v5

@@ I INTRODUCTION
Humans can anticipate the best viewpoints for a given object from only a single observation, thanks to their extensive experience and observation of various objects. This benefits from humans’ ability of spatial visualization and mental rotation [1, 2] (see the left half of Fig. LABEL:fig:Banner), which allows them to proactively adjust their position to gather more information. The viewpoint’s quality significantly impacts a robot’s perception and interaction with the surroundings. Poorly chosen viewpoints, often referred to as “accidental views”[3, 4], can obscure critical features of an object. In contrast, “canonical views” [5] provide perspectives that reveal the most informative features of an object, significantly enhancing the robot’s performance in tasks such as object detection[6, 7, 8], navigation[9, 10], 3D reconstruction [11, 12], and manipulation[13, 14]. Therefore, the ability to actively avoid accidental viewpoints and obtain better viewpoint is crucial for robots. This is known as active perception[15, 16], “where a robot purposefully selects what to perceive, how, and when to perceive it, driven by its current goals and interaction with the environment”.
@@ II Background & related works
The concept of “viewpoint” has been extensively discussed in computer graphics and virtual reality. The ability to select an optimal viewpoint is essential for accurately perceiving and interacting with objects in downstream tasks [17, 18, 19]. In this section, we begin with the methods used to evaluate viewpoint quality and then discuss the challenges of selecting and optimizing viewpoints.
@@ II-A Viewpoint Quality Evaluation
Currently, we do not have a unified definition of a ”good viewpoint” [22, 23]. Since Palmer et al. first introduced the concept of a ”Canonical Perspective” [24] in 1981, a wealth of work has emerged on how to evaluate viewpoints and select the best viewpoint. Various existing viewpoint theories have evaluated viewpoints based on different criteria, such as area [25, 26], silhouette [27, 28], depth [29], surface curvature [30, 31], semantics [32, 33], aesthetics[34], etc.